a chemist has two solutions of HCL. One has a 40% concentration and the other has a 25% concentration. how many liters of each solution must be mixed to obtain 65liters of a 29% solution
Answer by mananth(16946)

percent percentage ---------------- quantity
Acid I 40 ---------------- x gal
Acid II 25 ---------------- 65 - x gal
Mixture 29.00% ---------------- 65

40 x + 25 ( 65 - x ) = 29 * 65

40 x + 1625 - 25 x = 1885
40 x - 25 x = 1885 - 1625
15 x = 260
/ 15
x = 17.33 gal 40.00% Acid I
47.67 gal 25.00% Acid II